OpenStudy (anonymous):

A certain radioactive element decays according to N=No e ^-kt. If 50% of the original element sample remains after 9 days, what is the decay rate?

OpenStudy (anonymous):

you know if \(t=9\) you have half of what you started with set \[e^{9k}=.5\] and solve for \(k\)

OpenStudy (anonymous):

you got this? solve in two steps only

OpenStudy (anonymous):

how do you solve for k?

OpenStudy (anonymous):

1) write in equivalent logarithmic form (take the log) 2) divide by 9

OpenStudy (anonymous):

\[e^{9k}=.5\iff 9k=\ln(.5)\iff k=\frac{\ln(.5)}{9}\]

OpenStudy (anonymous):

then if you want a decimal, use a calculator

OpenStudy (anonymous):

i have no idea what that means haha. so is it: k= .1831?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait is ln the button on the calculator for e?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=ln%28.5%29%2F9 usually \(ln\) is the button and \(e\) requires the shift key

OpenStudy (anonymous):

ok, so i have to use the shift key right? i tried doing e ^.5 which was 1.64 and then divide that by 9 which was .1831

OpenStudy (anonymous):

oh, the link helped. sorry we didn't learn this yet so I have no idea what im talking about

OpenStudy (anonymous):

so the rate is k= -7.7 x 10 ^-2